6-3/4÷3=? & 8÷1-1/3=? please help me I'm having the hardest time trying to figure out these type of problems and fractions are my weakness. i've read many fractions problems but i still dan't get it.
Found 2 solutions by nycsub_teacher, KnightOwlTutor:
Answer by nycsub_teacher(90)

I assume that the fractions are mixed numbers as stated.
6-3/4 ÷ 3 =
I would change 6_3/4 to an improper fraction as step one.
So, 6_3/4 = 27/4
We now have:
27/4 divided by 3
You can write any number over 1 because any number divided by 1 = itself.
In other words, 3 can be written as 3/1.
27/4 divided by 3/1
We reverse 3/1 to 1/3 and change the problem from division to multiplication.
We do this because multiplication and division are inverses of each other.
We now have 27/4 X 1/3 = 9/4
Do the same with the other question.
Understand?




Answer by KnightOwlTutor(293)

first you solve for 6-3/4 You need to find the common denominator for both numbers in this case it is four. the number 6 becomes 24/4 you haven't changed the number it is still the same quantity but you can now perform the subtraction part with the fraction.64/4-3/4=21/4 divided by 3=7/4
8 divided by 1 is still eight but we need to find a common denominator to solve the problem in this case it is 3. 24/3-1/3=23/3